A 0. 750 g sample of gas PX3 is in a sealed 542 mL bulb at 26 and 1.00 atm of pressure. identify element X and name the gas.


1
Expert's answer
2021-01-11T03:53:06-0500

mass of sample = 0.750g

V = 542mL = 542 × 10-6 m3

T = 26°C = 273K + 26 = 299K

P = 1.00atm = 101325Nm-2


Using the Ideal gas law,

PV = nRT

101325 × 542 × 10-6 = n × 8.314 × 299

n = "\dfrac{101325×542×10^{-6}}{8.314×299} = 0.0221"


Therefore the sample of gas with a mass 0.750g contains 0.0223 moles of the compound


0.0221 moles = 0.750g

1 mole = x


x = "\dfrac{0.750}{0.0221}"g/mol = 33.94g/mol



PX3 = 33.94g/mol

30.97g/mol + (3 × X) = 33.94g/mol

3x = 2.97g/mol

x = 0.99g/mol


The only element with that small a molar mass is Hydrogen. Therefore, x = Hydrogen


Since x = H, the chemical formula for the chemical compound is PH3 which is another name for Phosphine with an IUPAC name Phosphane.